# Translating predictive distributions into informative priors

### v3

@@ Abstract
When complex Bayesian models exhibit implausible behaviour, one solution is to assemble available information into an informative prior. Challenges arise as prior information is often only available for the observable quantity, or some model-derived marginal quantity, rather than directly pertaining to the natural parameters in our model. We propose a method for translating available prior information, in the form of an elicited distribution for the observable or model-derived marginal quantity, into an informative joint prior. Our approach proceeds given a parametric class of prior distributions with as yet undetermined hyperparameters, and minimises the difference between the supplied elicited distribution and corresponding prior predictive distribution. We employ a global, multi-stage Bayesian optimisation procedure to locate optimal values for the hyperparameters. Three examples illustrate our approach: a cure-fraction survival model, where censoring implies that the observable quantity is a priori a mixed discrete/continuous quantity; a setting in which prior information pertains to $R^{2}$ – a model-derived quantity; and a nonlinear regression model.
@@ 2 Methodology
Uniqueness   Multiple equally faithful prior distributions may exist in complex models, meaning we must distinguish between such priors based on properties other than just faithfulness if a unique prior is desired. One approach is to choose the prior with the largest marginal standard deviations (see Section 2.3), but other properties could be used similarly. The challenge of uniqueness has been noted by Stefan et al. (2022).
@@
2.1 The target predictive distribution $\text{T}(Y)$
We assume a target predictive distribution, with cumulative distribution function (CDF) $\text{T}(Y)$, for the observable quantity $Y\in\mathcal{Y}\subseteq{}\mathbb{R}$ has been chosen. We recommend chosing this using predictive elicitation (Kadane & Wolfson, 1998), in which experts are queried about the observable quantity at a small number quantiles, then an appropriate parametric distribution is fitted to the elicited values (O’Hagan et al., 2006, chap. 6). However, our approach applies regardless of how $\text{T}(Y)$ is selected, although we assume it can be described by a (mixture of) standard distributions and that samples can be drawn from it.
We describe our methodology in the more general setting in which we wish to elicit information about the observable quantity $Y$ conditional on some known values of a covariate $X\in\mathcal{X}\subseteq\mathbb{R}^{C}$. For example, when using the linear model $Y=X\beta+\varepsilon$ we may elicit information about $Y$ at a fixed set of values for $X$ corresponding to an experimental design. Specifically we suppose the target CDF $\text{T}(Y\mid X_{r})$ has been elicited at $R$ values of the covariate vector denoted $\{X_{r}\}_{r=1}^{R}$, which we stack in the matrix $\boldsymbol{X}=\left[X_{1}^{\top}\cdots X_{R}^{\top}\right]\in\boldsymbol{\mathcal{X}}\subseteq\mathbb{R}^{R\times C}$.
@@ 2.2 Predictive discrepancy (primary objective)
Consider a joint model for observables $Y$ and parameters $\theta\in\Theta\subseteq\mathbb{R}^{Q}$, given hyperparameters $\lambda\in\Lambda\subset\mathbb{R}^{L}$ and covariates $X$. This joint model has CDF $\text{P}(Y,\theta\mid\lambda,X)$ and prior predictive CDF $\text{P}(Y\mid\lambda,X)$ for $Y$. We assume that each target $\text{T}(Y\mid X_{r})$ has identical support to $\text{P}(Y\mid\lambda,X_{r})$. We denote $\text{T}(Y\mid\boldsymbol{X})=\prod_{r=1}^{R}\text{T}(Y\mid X_{r})$, with $\text{P}(Y\mid\lambda,\boldsymbol{X})$ and $\text{P}(\theta\mid\lambda,\boldsymbol{X})$ defined analogously.
@@
Inspired by the Cramér-von Mises (von Mises, 1947) and Anderson-Darling (Anderson & Darling, 1952) distributional tests we define, for CDFs $\text{M}(Y)$ and $\text{P}(Y)$, two options for the discrepancy function,
@@
Both discrepancies are proper scoring rules (Gneiting & Raftery, 2007) as they are minimised iff $\text{M}(Y)=\text{P}(Y)$ for all $Y\in\mathcal{Y}$. Supposing $\text{P}(Y\mid\lambda,X_{r})$ is flexible enough to exactly match $\text{T}(Y\mid X_{r})$ for some $\lambda^{*}$, then both discrepancies will yield the same $\lambda^{*}$. Differences arise when $\text{P}(Y\mid\lambda,X_{r})$ is insufficiently flexible, in which case the Anderson-Darling (AD) discrepancy $d^{\text{AD}}$ places more emphasis on matching the tails of two CDFs.
@@ 2.3 Regularising estimates of $\lambda^{*}$ (secondary objective)
There often are many optimal values $\lambda^{*}$ that yield values of $\tilde{D}(\lambda^{*}\mid\boldsymbol{X})$ that are practically indistinguishable (noted by da Silva et al., 2019) but with immensly differing prior distributions $\text{P}(\theta\mid\lambda^{*},\boldsymbol{X})$. That is, there are many equally faithful priors. This is not surprising because we are providing information only on $Y$, which is typically of lower dimension than $\theta$. A particularly challenging case for uniqueness is in models with additive noise forms, such as (12); in this case it will generally be necessary to fix a prior for the variance using knowledge of the measurement process.
To handle more general cases of lack of uniqueness, we seek to encode the following principle into our methodology: given two estimates of $\lambda^{*}$ which have equivalent values of $\tilde{D}(\lambda^{*}\mid\boldsymbol{X})$, we prefer the one with the larger variance for $\text{P}(\theta\mid\lambda^{*},\boldsymbol{X})$. We accomplish this by defining a secondary objective $\tilde{N}(\lambda\mid\boldsymbol{X})$ formed from a suitable function $n(\theta)$ with
@@
In this paper we consider only one form for this: the (negative) mean of the marginal log standard deviations of each of the $Q$ components of $\theta\in\Theta\subseteq\mathbb{R}^{Q}$.
@@ 2.4 Algorithm and optimisation
This optimum depends on $\kappa$, which will usually be difficult to assess. However, using multi-objective optimisation we can evaluate (5) for any $\kappa$ without needing to redo the optimsation step, and thus plot Pareto frontiers for a wide range of values $\kappa\in\mathcal{K}$ coloured by loss, with the minimum loss point indicated. These can guide our choice of $\kappa$: we can seek a value of $\kappa$ with the minimum loss point not on the extreme of the Pareto frontier, since we would like to balance the two objectives. This approach is particularly useful in settings where the scales of the two optima differ markedly, which we further discuss in Supplement S2. Where it is feasible to replicate the optimisation procedure, we can additionally seek a choice of $\kappa$ that leads to Pareto frontiers with minimal variability across replicates, since this suggests the optimal solution can be estimated reliably.
@@
In stage one we minimise $\tilde{D}(\lambda\mid\boldsymbol{X})$ using controlled random search 2 (CRS2) with local mutation (Kaelo & Ali, 2006), which we run for $N_{\text{CRS2}}$ iterations. We make use of the final optimum value $\lambda^{*}$, as well as the $N_{\text{CRS2}}$ trial points to obtain a design $\mathcal{D}$ for the next stage. The design comprises values of $\lambda$, and their corresponding values of $\log(\tilde{D}(\lambda\mid\boldsymbol{X}))$. A (small) number of padding points $N_{\text{pad}}$ are added to $\mathcal{D}$ for numerical robustness in stage 2. The result is the design $\mathcal{D}=\left\{\lambda_{i},\log(\tilde{D}(\lambda_{i}\mid\boldsymbol{X}))\right\}_{i=1}^{N_{\text{design}}+N_{\text{pad}}}$. Whilst CRS2 was not designed to minimise noisy functions, it appears empirically robust to small quantities of noise.
@@ 3.1 Calibrating a cure fraction survival model
Cure models (Amico & Van Keilegom, 2018) for survival data are useful when a cure mechanism is physically plausible a priori, and when individuals are followed up for long enough to be certain all censored individuals in our data are “cured”. Such lengthy follow ups are not always possible, but a cure model remains plausible when a large fraction of the censored observations occur after the last observed event time. However, we cannot distinguish in the right tail of the survival time distribution between censored uncured individuals and genuinely cured individuals. We suppose here that we possess prior knowledge on the fraction of individuals likely to be cured, and the distribution of event times amongst the uncured, and seek to translate this information into a prior. This setting is interesting because the target distribution is of mixed discrete/continuous type, due to censoring. Additionally, we specify a model with a nontrivial correlation structure, about which we wish to specify an informative prior, which is known to be challenging.
@@ 3.1.1 Model
We will seek to identify optimal values of the hyperparamers $\lambda=(\alpha,\beta,\mu_{0},\allowbreak\sigma^{2}_{0},s_{\beta},\boldsymbol{\omega},\allowbreak\boldsymbol{\eta},\allowbreak a_{\pi},b_{\pi})^{\top}$, with $\pi\sim\text{Beta}(a_{\pi},b_{\pi})$, $\gamma\sim\text{Gamma}(\alpha,\beta)$, $\beta_{0}\sim\text{Normal}(\mu_{0},\sigma_{0}^{2})$ and $\boldsymbol{\beta}\sim\allowbreak\text{MVSkewNormal}(\boldsymbol{0},\allowbreak\boldsymbol{S},\boldsymbol{\eta})$, with $\boldsymbol{S}=\text{diag}(s_{\beta})\,\,\boldsymbol{\Omega}\,\,\text{diag}(s_{\beta})$ where $s_{\beta}$ is the prior marginal scale of $\boldsymbol{\beta}$ and $\boldsymbol{\Omega}$ is parameterised by $\boldsymbol{\omega}=(\omega_{1},\ldots,\omega_{6})^{\top}\in[-1,1]^{6}$ that uniquely determine its Cholesky factor. The skewness is necessary to incorporate the nonlinear relationship between the hazard and the effect of the covariates, and a covariance structure is used to account for fact that not all the elements of $\boldsymbol{\beta}$ can be large simultaneously. Further details are in Supplement S4.
@@ 3.1.2 Results
We use $S_{r}=10^{4}$ prior predictive samples and $I_{r}=5\times 10^{3}$ importance samples to estimate the discrepancy; and optimise using $N_{\text{CRS2}}=2000$ CRS2 iterations, followed by $N_{\text{batch}}=3$ batches of Bayesian optimisation with $N_{\text{BO}}=200$ iterations per batch, carrying $N_{\text{design}}=60$ points between batches. We select $\kappa$ by inspecting the Pareto frontiers for $\kappa\in\{0.1,0.2,0.3,0.5,1,2\}$ (Figure 2 and Supplement S4.2). Except for the maximum and minimum values, which yield minimum loss points on the extremes of the Pareto frontier, the minimum loss point is insensitive to a wide range of $\kappa$ values. We select $\kappa=0.3$ which simultaneously minimises variability in loss and both objectives.
The values of the loss and discrepancy functions at $\lambda^{*}$ across replicas are tightly distributed (see Supplement S4.3), which indicates replicability. We assess faithfulness by inspecting Figure 3. Across all replicates and discrepancies the estimated optimal prior predictive distribution is highly faithful to the target.
Figure 3: Estimated optimal prior predictive densities $\text{p}(Y_{n}\mid\lambda^{*})$ (red/blue lines and dots) and target densities $\text{t}(Y_{n}\mid C_{N})$ (black lines and crosses) for individual $n=9$ (other individuals are visually indistinguishable).
Figure 4: Estimated optimal prior marginal densities of $\text{p}(\theta\mid\lambda^{*})$ for each component of $\theta$ ($\beta_{2},\dots,\beta_{4}$, not shown, are near identical to $\beta_{1}$). Both axes are manually truncated for readability.
Figure 4 displays the marginals of $\theta$ for each independent replicate. The single objective approach (i.e. only optimising predictive discrepancy, shown in blue) consistently locates the degenerate, non-unique solution where all the variation in the uncensored event times is attributed to the baseline hazard shape $\gamma$ and the intercept $\beta_{0}$: i.e. all the mass for $\boldsymbol{\beta}$ (the regression coefficients) is close to 0. The combination of $\gamma$ and $\beta_{0}$ is evidently poorly identified, and further calculation reveals that only the derived product $\gamma\exp{\beta_{0}}$ is uniquely determined. Given the inter-replicate consistency previously observed in Supplement S4.3 we infer that the solution is not unique. In the multi-objective approach, there is a preference for the optima surrounding $\gamma\approx 7.5,\beta_{0}\approx-10$; an improvement in uniqueness over the single objective approach, but imperfect.

### v4

@@ Abstract
“When complex Bayesian models exhibit implausible behaviour, one solution is to assemble available information into an informative prior. Challenges arise as prior information is often only available for the observable quantity, or some model-derived marginal quantity, rather than directly pertaining to the (usually latent) parameters in our model. We propose a method for translating available prior information, in the form of an elicited distribution for the observable or model-derived marginal quantity, into an informative joint prior. Our approach proceeds given a parametric class of prior distributions with as yet undetermined hyperparameters, and minimises the difference between the supplied elicited distribution and corresponding prior predictive distribution. We employ a global, multi-stage Bayesian optimisation procedure to locate optimal values for the hyperparameters. Three examples illustrate our approach: a cure-fraction survival model, where censoring implies that the observable quantity is a priori a mixed discrete/continuous quantity; a setting in which prior information pertains to $R^{2}$ – a model-derived quantity; and a nonlinear regression model.”
@@ 2 Methodology
Uniqueness   Multiple equally faithful prior distributions may exist in complex models, meaning we must distinguish between such priors based on properties other than just faithfulness if a unique prior is desired. For example, if maximising prior uncertainty whilst retaining faithfulness is desired, then we could choose the prior with the largest marginal standard deviations (see Section 2.3). Other properties could be used similarly. The challenge of uniqueness has been noted by Stefan et al. (2022).
@@
2.1 Setup
Consider a joint probability distribution for an observable $Y\in\mathcal{Y}\subseteq{}\mathbb{R}$ and parameters $\theta\in\Theta\subseteq\mathbb{R}^{Q}$, given hyperparameters $\lambda\in\Lambda\subset\mathbb{R}^{L}$. This distribution has cumulative distribution function (CDF) $\text{P}(Y,\theta\mid\lambda)$ and prior predictive CDF $\text{P}(Y\mid\lambda)$ for $Y$. We suppose a target predictive distribution, with CDF $\text{T}(Y)$, for the observable quantity $Y$ has been chosen, and that this encapsulates our prior knowledge about $Y$. Our primary aim is to choose $\lambda$ so that the prior predictive $\text{P}(Y\mid\lambda)$ is faithful to the target $\text{T}(Y)$.
We assume that the target $\text{T}(Y)$ can be described by a (mixture of) standard distributions and that samples can be drawn from it; but we do not require $\text{T}(Y)$ to be in the same parametric family as the prior predictive $\text{P}(Y\mid\lambda)$, since this is often unavailable in closed-form. We recommend choosing $\text{T}(Y)$ using predictive elicitation (Kadane & Wolfson, 1998), in which an appropriate parametric distribution is fitted to a small number of quantiles (of the observable quantity) elicited from experts (O’Hagan et al., 2006, chap. 6).
We describe our methodology in a slightly more general setting in which the observable quantity $Y$ is conditional on a covariate $X\in\mathcal{X}\subseteq\mathbb{R}^{C}$, with joint probability distribution CDF $\text{P}(Y,\theta\mid\lambda,X)$ and prior predictive CDF $\text{P}(Y\mid\lambda,X)$. We assume information has been elicited about $Y$ at a fixed set of values for $X$. Specifically we suppose the target CDF $\text{T}(Y\mid X_{r})$ has been elicited at $R$ values of the covariate vector denoted $\{X_{r}\}_{r=1}^{R}$, which we stack in the matrix $\boldsymbol{X}=\left[X_{1}^{\top}\cdots X_{R}^{\top}\right]\in\boldsymbol{\mathcal{X}}\subseteq\mathbb{R}^{R\times C}$. We assume that each target $\text{T}(Y\mid X_{r})$ has identical support to $\text{P}(Y\mid\lambda,X_{r})$. We denote $\text{T}(Y\mid\boldsymbol{X})=\prod_{r=1}^{R}\text{T}(Y\mid X_{r})$, with $\text{P}(Y\mid\lambda,\boldsymbol{X})$ and $\text{P}(\theta\mid\lambda,\boldsymbol{X})$ defined analogously.
@@ 2.2 Predictive discrepancy (primary objective)
Many forms of discrepancy function $d(\cdot,\cdot)$ could be adopted, but restricting to proper scoring rules (Gneiting & Raftery, 2007), which are minimised iff $\text{P}(Y\mid\lambda,X_{r})=\text{T}(Y\mid\lambda)$ for all $Y\in\mathcal{Y}$, is intuitive. In this case, if $\text{P}(Y\mid\lambda,X_{r})$ is flexible enough to exactly match $\text{T}(Y\mid X_{r})$ for some $\lambda^{*}$, then any such discrepancy will yield the same $\lambda^{*}$. Differences arise when $\text{P}(Y\mid\lambda,X_{r})$ is insufficiently flexible, with discrepancy functions differing in placement of emphasis.
CDF-based discrepancies are appealing because they are widely-applicable, so inspired by the Cramér-von Mises (von Mises, 1947) and Anderson-Darling (Anderson & Darling, 1952) distributional tests we define, for CDFs $\text{M}(Y)$ and $\text{P}(Y)$:
@@
The Anderson-Darling (AD) discrepancy $d^{\text{AD}}$ places more emphasis than Cramér-von Mises (CvM) on matching the tails of two CDFs.
Another option is either direction of Kullback-Leibler (KL) divergence,
$d^{\text{KL-fwd}}=\text{KL}(\text{M}(Y)\;\|\;\text{P}(Y)),\quad d^{\text{KL-rev}}=\text{KL}(\text{P}(Y)\;\|\;\text{M}(Y)).$ (3)
We consider the form of KL divergences detailed in Supplement S4.
@@ 2.3 Regularising estimates of $\lambda^{*}$ (secondary objective)
There often are many optimal values $\lambda^{*}$ that yield values of $\tilde{D}(\lambda^{*}\mid\boldsymbol{X})$ that are practically indistinguishable (noted by da Silva et al., 2019) but with immensly differing prior distributions $\text{P}(\theta\mid\lambda^{*},\boldsymbol{X})$. That is, there are many equally faithful priors. This is not surprising because we are providing information only on $Y$, which is typically of lower dimension than $\theta$. A particularly challenging case for uniqueness is in models with additive noise forms, such as (13); in this case it will generally be necessary to fix a prior for the noise using knowledge of the measurement process.
To handle more general cases of lack of uniqueness, we define a secondary objective $\tilde{N}(\lambda\mid\boldsymbol{X})$, typically via a function $n(\theta)$ with
@@
This objective can be chosen by practitioners to promote or inhibit properties of the prior predictive $\text{P}(Y\mid\lambda,X_{r})$ as desired.
As an example, we demonstrate how to encode a preference for maximising prior uncertainty, as is commonly desired in the absence of contrary prior knowledge. Specifically, given two estimates of $\lambda^{*}$ which have equivalent values of $\tilde{D}(\lambda^{*}\mid\boldsymbol{X})$, we prefer the one with the larger variance for $\text{P}(\theta\mid\lambda^{*},\boldsymbol{X})$. This preference could be encoded in several ways: a simple option is the (negative) mean of the marginal log standard deviations across the $Q$ components of $\theta\in\Theta\subseteq\mathbb{R}^{Q}$.
@@ 2.4 Algorithm and optimisation
This optimum depends on $\kappa$, which will usually be difficult to assess. However, using multi-objective optimisation we can evaluate (6) for any $\kappa$ without needing to redo the optimisation step, and thus plot Pareto frontiers for a wide range of values $\kappa\in\mathcal{K}$ coloured by loss, with the minimum loss point indicated. These can guide our choice of $\kappa$: we can seek a value of $\kappa$ with the minimum loss point not on the extreme of the Pareto frontier, since we would like to balance the two objectives. This approach is particularly useful in settings where the scales of the two optima differ markedly, which we further discuss in Supplement S2. Where it is feasible to replicate the optimisation procedure, we can additionally seek a choice of $\kappa$ that leads to Pareto frontiers with minimal variability across replicates, since this suggests the optimal solution can be estimated reliably.
@@
In stage one we minimise $\tilde{D}(\lambda\mid\boldsymbol{X})$ using controlled random search 2 (CRS2) with local mutation (Kaelo & Ali, 2006), which we run for $N_{\text{CRS2}}$ iterations. We use the final optimum value $\lambda^{*}$, as well as the $N_{\text{CRS2}}$ trial points, to obtain a design $\mathcal{D}$ for the next stage. The design comprises values of $\lambda$, and their corresponding values of $\log(\tilde{D}(\lambda\mid\boldsymbol{X}))$. A (small) number of padding points $N_{\text{pad}}$ are added to $\mathcal{D}$ for numerical robustness in stage 2. The result is the design $\mathcal{D}=\left\{\lambda_{i},\log(\tilde{D}(\lambda_{i}\mid\boldsymbol{X}))\right\}_{i=1}^{N_{\text{design}}+N_{\text{pad}}}$. Whilst CRS2 was not designed to minimise noisy functions, empirically it appears robust to small quantities of noise.
@@ 3.1 Calibrating a cure fraction survival model
Cure models (Amico & Van Keilegom, 2018) for survival data are useful when a cure mechanism is physically plausible a priori, and when individuals are followed up for long enough to be certain all censored individuals in our data are “cured”. Such lengthy follow ups are not always possible, but a cure model remains plausible when a large fraction of the censored observations occur after the last observed event time. However, we cannot distinguish in the right tail of the survival time distribution between censored uncured individuals and genuinely cured individuals. We suppose here that we possess prior knowledge on the fraction of individuals likely to be cured, and the distribution of event times amongst the uncured, and seek to translate this information into a prior. We consider the CDF-based CvM and AD discrepancies in this example because the target distribution is of mixed discrete/continuous type (due to censoring). Additionally, we specify a model with a nontrivial correlation structure, about which we wish to specify an informative prior, which is known to be challenging.
@@ 3.1.1 Model
We will seek to identify optimal values of the hyperparamers $\lambda=(\alpha,\beta,\mu_{0},\allowbreak\sigma^{2}_{0},s_{\beta},\boldsymbol{\omega},\allowbreak\boldsymbol{\eta},\allowbreak a_{\pi},b_{\pi})^{\top}$, with $\pi\sim\text{Beta}(a_{\pi},b_{\pi})$, $\gamma\sim\text{Gamma}(\alpha,\beta)$, $\beta_{0}\sim\text{Normal}(\mu_{0},\sigma_{0}^{2})$ and $\boldsymbol{\beta}\sim\allowbreak\text{MVSkewNormal}(\boldsymbol{0},\allowbreak\boldsymbol{S},\boldsymbol{\eta})$, with $\boldsymbol{S}=\text{diag}(s_{\beta})\,\,\boldsymbol{\Omega}\,\,\text{diag}(s_{\beta})$ where $s_{\beta}$ is the prior marginal scale of $\boldsymbol{\beta}$ and $\boldsymbol{\Omega}$ is parameterised by $\boldsymbol{\omega}=(\omega_{1},\ldots,\omega_{6})^{\top}\in[-1,1]^{6}$ that uniquely determine its Cholesky factor. The skewness is necessary to incorporate the nonlinear relationship between the hazard and the effect of the covariates, and a covariance structure is used to account for fact that not all the elements of $\boldsymbol{\beta}$ can be large simultaneously. Further details are in Supplement S5.
@@ 3.1.2 Results
We use $S_{r}=10^{4}$ prior predictive samples and $I_{r}=5\times 10^{3}$ importance samples to estimate the discrepancy; and optimise using $N_{\text{CRS2}}=2000$ CRS2 iterations, followed by $N_{\text{batch}}=3$ batches of Bayesian optimisation with $N_{\text{BO}}=200$ iterations per batch, carrying $N_{\text{design}}=60$ points between batches. We select $\kappa$ by inspecting the Pareto frontiers for $\kappa\in\{0.1,0.2,0.3,0.5,1,2\}$ (Figure 2 and Supplement S5.2). Except for the maximum and minimum values, which yield minimum loss points on the extremes of the Pareto frontier, the minimum loss point is insensitive to a wide range of $\kappa$ values. We select $\kappa=0.3$ which simultaneously minimises variability in loss and both objectives.
The values of the loss and discrepancy functions at $\lambda^{*}$ across replicas are tightly distributed (see Supplement S5.3), which indicates replicability. Across all replicates and discrepancies the estimated optimal prior predictive distribution is highly faithful to the target, as illustrated for individual $n=9$ in Figure 3 (other individuals are visually indistinguishable).
Figure 3: Estimated optimal prior predictive densities $\text{p}(Y_{n}\mid\lambda^{*})$ (red/blue lines and dots) and target densities $\text{t}(Y_{n}\mid C_{N})$ (black lines and crosses).
Figure 4: Estimated optimal prior marginal densities of $\text{p}(\theta\mid\lambda^{*})$ for each component of $\theta$ ($\beta_{2},\dots,\beta_{4}$, not shown, are near identical to $\beta_{1}$). Both axes are truncated for readability.
Figure 4 displays the marginals of $\theta$ for each independent replicate. The single objective approach (i.e. only optimising predictive discrepancy, shown in blue) consistently locates the degenerate, non-unique solution where all the variation in the uncensored event times is attributed to the baseline hazard shape $\gamma$ and the intercept $\beta_{0}$: i.e. all the mass for $\boldsymbol{\beta}$ (the regression coefficients) is close to 0. The combination of $\gamma$ and $\beta_{0}$ is far from unique, and further calculation reveals that only the derived product $\gamma\exp{\beta_{0}}$ is uniquely determined. Given the inter-replicate consistency previously observed in Supplement S5.3 we infer that the solution is not unique. In the multi-objective approach, there is a preference for the optima surrounding $\gamma\approx 7.5,\beta_{0}\approx-10$; an improvement in uniqueness over the single objective approach, but imperfect.